a.      Calculate the pH of a solution consisting of 1.35M acetic acid (CH3COOH)

 (pKa= 1.8 x 10-5) and 2.83 M sodium acetate (NaC2H3O2).


1
Expert's answer
2022-02-28T11:04:58-0500

pH= pKa + log([base]÷[acid])

=-log(1.8×10-5 ) + log (2.83÷1.35)

= 4.74+0.321

= 5.06